# Improved real-space parallelizable matrix-product state compression and its application to unitary quantum dynamics simulation

### v1

@@ III.5 Improved parallel MPS compression (IPMC)
Inspired by the new insights from Theorems 2 and 3, we propose the IPMC method: starting from an MPS in the Vidal form, which is usually the resulting MPS of the previous TEBD evolution and has the bond dimension $d\chi$ ($d$ is the dimension of the local Hilbert space, i.e., $d=2$ for the qubit case), we perform a parallel MPS compression process (discussed in Sec. III.2) followed by a norm stabilization step (discussed in Sec. III.3.2) to compress the bond dimension back to $\chi$. Then, we perform $g$ PtSU steps to partially recover the canonical form. This parallel MPS compression method is fully real-space parallelizable, requiring only communication between neighboring sites [also see Fig. 2(d)]. In the next section, we will apply this method to simulate unitary quantum dynamics and examine its efficiency through extensive numerical experiments.
@@ IV Application to the simulation of unitary quantum dynamics
where $\hat{U}_{i}^{d}$ is the $i$th local unitary operator in the $d$th layer, $m(d)$ is the number of local unitary operators in the $d$th layer, and $D$ is the total number of layers, i.e., depth, of the quantum circuit [26].
The above decomposition is, in general, approximate and, moreover, it is not unique. Therefore, for the convenience of applying MPS-based algorithms, one usually constructs $\{\hat{U}_{i}^{d}\}$ by only involving adjacent qubits in a 1D path going through all the qubits in the system [see Fig. 2(b) as an example and also Appendix B]. Therefore, in the rest of this section, we discuss the simulations of unitary quantum dynamics by considering their 1D quantum circuit representations.
@@ IV.1 Parallel time-evolving block decimation (pTEBD) algorithm
After completing the simulation of all $D$ layers of gates in the quantum circuit, the resulting MPS approximates the final state of the unitary quantum dynamic process $|\Psi(t)\rangle$. We can then use standard MPS techniques to exactly evaluate the expectation values of observables [19] or perfectly sample bit strings from this state [38] to mimic quantum devices after restoring the canonical form of the obtained MPS representing $|\Psi(t)\rangle$. The flow chat of the pTEBD algorithm is provided in Algorithm 1.
Figure 3: Wavefunction fidelity $\mathcal{F}$ as a function of physical circuit depth $D$ (refer to Appendix B) obtained using the sequential MPS algorithm and the pTEBD algorithm with a fixed MPS bond dimension $\chi$: (a) RQC-1D with $N=25$, (b) PQC-1D with $N=24$, (c) RQC-2D with $L_{x}=5$ and $L_{y}=5$ (i.e., $N=25$), and (d) PQC-2D with $L_{x}=4$ and $L_{y}=6$ (i.e., $N=24$). The dashed lines (crosses) represent the average fidelity ¯$\bar{\mathcal{F}}_{\rm MPS(pTEBD)}$ over 10 simulations of the same quantum circuit but with different sets of random parameters obtained using the sequential MPS (pTEBD) algorithm. The shades (bars) indicate the minimal and maximal fidelities among these 10 simulations using the sequential MPS (pTEBD) algorithm. The color intensities of crosses and bars represent the pTEBD results with $g=0,1,\text{and}~{}2$ (from light to dark).
@@ IV.2 Accuracy, numerical stability, and performance of pTEBD
To comprehensively benchmark the accuracy, numerical stability, and performance of the pTEBD algorithm, we perform intensive simulations of typical 1D and 2D quantum circuits and compare the results with those obtained by sequential MPS simulations. For this purpose, we consider both random quantum circuits (RQCs) on 1D and 2D qubit arrays (RQC-1D and RQC-2D, respectively) and parametrized quantum circuits (PQCs), which are widely adopted in the variational quantum eigensolver (VQE) [39], on 1D and 2D qubit arrays (PQC-1D and PQC-2D, respectively). These four types of quantum circuits cover representative cases, providing concrete demonstrations for the practical feasibility of the pTEBD algorithm. The details of each circuit are described in Appendix B. Note also that in the sequential algorithm [7], an MPS is always in the canonical form and normalized to one, which is generally not the case in the pTEBD algorithm.
Figure 4: Wavefunction norm of the MPS as a function of physical circuit depth $D$ (refer to Appendix B) obtained using the pTEBD algorithm with a fixed MPS bond dimension $\chi$: (a) RQC-1D with $N=101$, (b) PQC-1D with $N=100$, (c) RQC-2D with $L_{x}=L_{y}=12$ (i.e., $N=144$), and (d) PQC-2D with $L_{x}=L_{y}=12$ (i.e., $N=144$). The open diamonds (squares) represent the averaged results of the wavefunction norm $n^{*}$ ($n$) over 10 simulations of the same quantum circuit but with different sets of random parameters, evaluated with (without) the wavefunction norm stabilization procedure. Black dashed lines indicate the norm equals to 1. Here we set $g=0$, i.e., no PtSU. Note that the results obtained using the pTEBD algorithm are almost completely overlapped to each other in this scale.
@@ IV.2.1 Accuracy
First of all, we assess the simulation accuracy of the pTEBD algorithm and demonstrate its ability to achieve precision comparable to that of the sequential MPS algorithm described in Ref. [7]. Here, the simulation precision is determined by the wavefunction fidelity
$\mathcal{F}=|\langle\Psi_{\rm exact}|\Psi(\chi)\rangle|~{},$ (37)
where $|\Psi_{\rm exact}\rangle$ represents the state obtained by the exact simulation and $|\Psi(\chi)\rangle$ is the state obtained using an MPS with the fixed bond dimension $\chi$. We perform the simulations for quantum circuits with up to 100 (60) physical circuit layers, i.e., before recompiling a quantum circuit to fit an MPS 1D path, in the 1D (2D) cases. The definitions of the physical and compiled circuit depth can be found in Appendix B. Each simulation is repeated 10 times with different sets of random parameters to evaluate the average fidelity, denoted as ¯$\bar{\mathcal{F}}_{\rm MPS(pTEBD)}$ in the sequential MPS (pTEBD) simulations. These results are summarized in Fig. 3.
The simulations for all these four types of quantum circuits show consistent results: for a given bond dimension $\chi$, the accuracy of the pTEBD simulation is comparable to the sequential MPS simulation. In most cases, we find that ¯$\bar{\mathcal{F}}_{\rm pTEBD}\approx\bar{\mathcal{F}}_{\rm MPS}$ and ¯$\bar{\mathcal{F}}_{\rm pTEBD}$ improves with an increase in the number $g$ of PtSU steps, while we observe that the PtSU steps have the opposite effect in some rare cases. Moreover, in cases with relatively small $\chi$ (for example, $\chi=16$ and 64), we notice that ¯$\bar{\mathcal{F}}_{\rm pTEBD}$ significantly increases with applying more PtSU steps. Since $\mathcal{F}$ is highly sensitive to the random parameters used in each quantum gate in the quantum circuit, we also shown its minimum and maximum among the 10 simulations with different sets of random parameters. These extremes obtained in the sequential MPS simulations and the pTEBD simulations are consistent with each other. In some cases [see Fig. 3(b)], the maximum value obtained in the pTEBD simulation can be larger than that obtained in the sequential MPS simulation. To more clearly present the fidelities obtained using these two algorithms, we also show the results for a quantum circuit with the same single set of random parameters in Appendix C.
@@
Figure 5: Elapsed time per circuit layer (averaged over 10 simulations with different sets of random parameters) versus the system size $N$ ($L_{x}=L_{y}$ in 2D cases) obtained using the sequential MPS algorithm (open circles) and the pTEBD algorithm (crosses) with a fixed MPS bond dimension $\chi$: (a) RQC-1D with 300 physical circuit layers, (b) PQC-1D with 300 physical circuit layers, (c) RQC-2D with 100 physical circuit layers, and (d) PQC-2D with 100 physical circuit layers. Note that the number of circuit layers in (c) and (d) is the total number of layers of the quantum circuit after recompiling it to fit an MPS 1D path, i.e., the compiled circuit depth (for more details, see Appendix B). We set $g=0$ for the pTEBD simulations.
@@ IV.2.2 Numerical stability
IV.2.3 Performance
Finally, let us assess the performance of the pTEBD algorithm from two perspectives. On the one hand, we examine the simulation elapsed time per circuit layer, which is defined as the total simulation elapsed time divided by compiled circuit depth (for more details, see Appendix B), versus the system size $N$, evaluating its performance in weak scaling. On the other hand, we compare the elapsed time required to achieve a given simulation precision, measured by wavefunction fidelity $\mathcal{F}$, in both sequential MPS and pTEBD simulations.
As illustrated in Fig. 5, for a fixed bond dimension $\chi$, the pTEBD algorithm consistently exhibits nearly constant simulation times for quantum circuits with various system sizes, achieving excellent weak scaling performance. This is attributed to the absence of any real-space sequential procedures and the lack of a need for global data communication in the pTEBD algorithm [also see Fig. 2(d)]. In contrast, the elapsed time per circuit layer in the sequential MPS algorithm increases linearly with the system size due to its unavoidable sweep procedure.
Figure 6: Wavefunction fidelity $\mathcal{F}$ versus elapsed time obtained using the sequential MPS algorithm (open circles) and the pTEBD algorithm (crosses) with a fixed physical circuit depth $D$ (refer to Appendix B): (a) RQC-1D with $N=25$, (b) PQC-1D with $N=24$, (c) RQC-2D with $L_{x}=L_{y}=5$ (i.e., $N=25$), and (d) PQC-2D with $L_{x}=4$ and $L_{y}=6$ (i.e., $N=24$). These results are averaged over 10 simulations of the same quantum circuit but with different sets of random parameters. We set $g=0$ for the pTEBD simulations.
It might be more insightful to compare the performance between these two algorithms by studying their elapsed time to achieve the same simulation precision, a measure similarly adopted in a recent study of a 2D tensor network algorithm [40]. This measure demonstrates the practical speedup of the pTEBD algorithm over its sequential counterpart. For this purpose, we reinterpret the fidelity results shown in Fig. 3 and replot these results as a function of the elapsed time of the simulation in Fig. 6. Remarkably, in all cases, the pTEBD simulation reaches the same $\mathcal{F}$ in polynomially shorter time (as also expected from Fig. 5) than the corresponding sequential MPS simulation, indicating its efficient utilization of the parallel computing environment.
@@ IV.2.3 Performance
In summary, we have proposed an improved parallel MPS compression method that can accurately compress the dimensions of all the virtual bonds in a constant time, regardless of the system size. Simultaneously, it stabilizes the wavefunction norm of the MPS, converging to a value around 1 that is bounded from both sides. Although both the accuracy and the norm stabilization are mathematically proved under the assumption of the canonical form, we have demonstrated its feasibility in simulating unitary quantum dynamics, where the canonical form slightly deviates globally. Moreover, we have numerically shown that the deviated canonical form resulting from the parallel truncation procedure can be gradually restored by appended PtSU steps. Additionally, we have provided a proof that a sufficient number of PtSU steps can drive any MPS to its canonical form.
Utilizing the IPMC method, we have proposed a fully real-space parallelizable pTEBD algorithm for efficiently simulating unitary quantum dynamics. Furthermore, we systematically benchmarked the pTEBD algorithm by simulating typical 1D and 2D quantum circuits. Our results demonstrate that the pTEBD algorithm achieves nearly perfect weak scaling performance even for very deep quantum circuits with hundreds of circuit layers, Additionally, it achieves the same simulation precision in polynomially shorter time compared to the sequential MPS algorithm.
@@ Acknowledgements.
We acknowledge the valuable discussions with Hong-Hao Tu, Kazuma Nagao, Hidehiko Kohshiro, and Shigetoshi Sota. Numerical simulations were performed using the high-performance tensor computing library GraceQ/tensor [41] on the Supercomputer Fugaku installed at RIKEN R-CCS (Project ID No. hp220217 and No. hp230293). This work is supported by Grant-in-Aid for Scientific Research (A) (No. JP21H04446) and Grant-in-Aid for Scientific Research (C) (No. JP22K03479) from MEXT, Japan, as well as JST COI-NEXT (Grant No. JPMJPF2221). Additionally, this work receives support from the Program for Promoting Research on the Supercomputer Fugaku (No. JPMXP1020230411) from MEXT, Japan, and from the COE research grant in computational science from Hyogo Prefecture and Kobe City through the Foundation for Computational Science.
Appendix A Approaching the canonical form through the application of PtSU
To measure how an MPS approach its canonical form with PtSU steps, we first define the distance to the canonical form as
where |||||$\cdot$||| | ⋅ | | represents the Frobenius norm of the matrix. This quantity is zero if and only if an MPS is in the canonical form. Notice that a similar quantity has been proposed in a recent study to examine the regauging of general tensor networks, where the nuclear norm is considered [42].
We investigate how $\mathcal{C}$ changes with the application of PtSU steps. To closely mimic the actual calculation task, we prepare the initial MPS as follows: We first construct a canonicalized random MPS with the bond dimension $\chi$. Each tensor element in this MPS is initialized to a complex number $z=a+ib$ with $a$ and $b$ chosen randomly from $[-1,1]$, and then its canonical form is constructed with the norm properly normalized to 1. Subsequently, we truncate all the virtual bonds to $\chi/2$, i.e., disregarding the $\chi/2$ smallest elements in each $\Lambda^{[i]}$, and renormalize the wavefunction norm to 1. We then perform PtSU steps on this MPS and calculate $\mathcal{C}$ after each PtSU step. The results are summarized in Fig. 7. Here, we fix $N=24$ and examine the cases of $\chi=32$, 128, 256 and 512.
Figure 7: Seim-log plot of $\mathcal{C}$ normalized by the initial value $\mathcal{C}_{0}$ as a function of PtSU steps. The initial MPS is prepared by parallelly truncating all the virtual bonds of a canonicalized random MPS to half of the bond dimension and manually normalizing the norm of the resulting MPS to 1. The results are averaged over 1000 initial MPSs with different sets of random parameters.
In Fig. 7, we observe that $\mathcal{C}$ rapidly and monotonically decreases with PtSU steps, especially at the beginning of several PtSU steps. Additionally, it more rapidly attains the canonical form in the case of a larger bond dimension. This implies that we can drive the MPS closer to the canonical form by appending a few PtSU steps after parallel truncation.
Appendix B Details of the quantum circuits adopted in the benchmarking simulations
To benchmark the pTEBD algorithm, we select random quantum circuits (RQCs) and parametrized quantum circuits (PQCs), the latter being commonly used in variational quantum algorithms, as the simulation tasks. RQCs have been widely employed in benchmarking NISQ devices [1, 43] and quantum computing simulators [15]. PQCs, being more featured circuits, can reflect the practical performance of devices and simulators for more realistic computational tasks. Therefore, these benchmarks allow for a thorough evaluation of the pTEBD algorithm’s performance in the most realistic and representative cases. In this Appendix, we provide explanations for the construction of these quantum circuits. More detailed features and applications of these quantum circuits can be found in the corresponding references cited below.
B.1 Physical circuit depth vs. compiled circuit depth

### v2

@@ III.5 Improved parallel MPS compression (IPMC)
Inspired by the new insights from Theorems 2 and 3, we propose the IPMC method: starting from an MPS in the Vidal form, which is usually the resulting MPS of the previous TEBD evolution and has the bond dimension $d_{\rm loc}\chi$ ($d_{\rm loc}$ is the dimension of the local Hilbert space, i.e., $d_{\rm loc}=2$ for the qubit case), we perform a parallel MPS compression process (discussed in Sec. III.2) followed by a norm stabilization step (discussed in Sec. III.3.2) to compress the bond dimension back to $\chi$. Then, we perform $g$ PtSU steps to partially recover the canonical form. This parallel MPS compression method is fully real-space parallelizable, requiring only communication between neighboring sites [also see Fig. 3(d)]. In the next section, we will apply this method to simulate unitary quantum dynamics and examine its efficiency through extensive numerical experiments.
@@ IV Application to the simulation of unitary quantum dynamics
where $\hat{U}_{i}^{d}$ is the $i$th local unitary operator in the $d$th layer, $m(d)$ is the number of local unitary operators in the $d$th layer, and $D$ is the total number of layers (i.e., depth) of the quantum circuit [32]. Notice that there exist other simulation approaches when $\hat{\mathcal{U}}(t)$ is associated with a Hamiltonian [41, 29].
Figure 3: Illustration of the pTEBD algorithm for a quantum circuit simulation. Starting from an MPS in (a), which is allowed to be slightly deviated from the canonical form, the state is evolved using the standard TEBD procedure but with all gates in the same layer treated parallelly in (b), and then truncate the state using the IPMC (see Sec. III.5) in (c) to update the MPS, which is subsequently used to simulate the quantum circuit with the next layer of gates. A possible realization of the distribution of data to each MPI process ${\rm P}_{l}$ and the data transfer among MPI processes in the pTEBD algorithm is schematically shown in (d). As multiple PtSU steps may be executed in the IPMC, the data transfer is consequently required several times, as indicated here simply by two arrows pointing forward and backward between neighboring MPI processes. Note also that single-qubit gates can be treated exactly without increasing the bond dimension by any MPS-based simulations, including the pTEBD algorithm.
The above decomposition is, in general, approximate and, moreover, it is not unique. Therefore, for the convenience of applying MPS-based algorithms, one usually constructs $\{\hat{U}_{i}^{d}\}$ only involving adjacent qubits in a 1D path going through all the qubits in the system [see Fig. 3(b) as an example and also Appendix A]. Therefore, in the rest of this section, we discuss the simulations of unitary quantum dynamics by considering their 1D quantum circuit representations.
@@ IV.1 Parallel time-evolving block decimation (pTEBD) algorithm
After completing the simulation of all $D$ layers of gates in the quantum circuit, the resulting MPS approximates the final state of the unitary quantum dynamic process $|\Psi(t)\rangle$. We can then use standard MPS techniques to exactly evaluate the expectation values of observables [23] or perfectly sample bit strings from this state [45] to mimic quantum devices after restoring the canonical form of the obtained MPS representing $|\Psi(t)\rangle$. The flow chart of the pTEBD algorithm is provided in Algorithm 1.
Figure 4: Wavefunction fidelity $\mathcal{F}$ as a function of physical circuit depth $D$ (for the definition, see Appendix A) obtained using the sequential MPS algorithm and the pTEBD algorithm with a fixed MPS bond dimension $\chi$: (a) RQC-1D with $N=25$, (b) PQC-1D with $N=24$, (c) RQC-2D with $L_{x}=5$ and $L_{y}=5$ (i.e., $N=25$), and (d) PQC-2D with $L_{x}=4$ and $L_{y}=6$ (i.e., $N=24$). The dashed lines (crosses) represent the average fidelity ¯$\bar{\mathcal{F}}_{\rm MPS(pTEBD)}$ over 10 simulations of the same quantum circuit but with different sets of random parameters obtained using the sequential MPS (pTEBD) algorithm. The shades (bars) indicate the minimal and maximal fidelities among these 10 simulations using the sequential MPS (pTEBD) algorithm. The color intensities of crosses and bars represent the pTEBD results with $g=0,1,\text{and}~{}2$ (from light to dark).
@@ IV.2 Accuracy, numerical stability, and performance of pTEBD
To comprehensively benchmark the accuracy, numerical stability, and performance of the pTEBD algorithm, we perform intensive simulations of typical 1D and 2D quantum circuits and compare the results with those obtained by sequential MPS simulations. For this purpose, we consider both random quantum circuits (RQCs) on 1D and 2D qubit arrays (RQC-1D and RQC-2D, respectively) and parametrized quantum circuits (PQCs), which are widely adopted in the variational quantum eigensolver (VQE) [46], on 1D and 2D qubit arrays (PQC-1D and PQC-2D, respectively). These four types of quantum circuits cover representative cases, providing concrete demonstrations for the practical feasibility of the pTEBD algorithm. The details of each circuit are described in Appendix A. Note also that in the sequential algorithm described in Ref. [7], which we employ as the benchmark algorithm, referred to as the sequential MPS algorithm, in the following numerical experiments, an MPS is always in the canonical form and normalized to one, which is generally not the case in the pTEBD algorithm. We also provide an additional benchmark test of the pTEBD algorithm applied to simulate the quantum Fourier transformation algorithm in Appendix C.
Figure 5: Wavefunction norm of the MPS as a function of physical circuit depth $D$ (for the definition, see Appendix A) obtained using the pTEBD algorithm with a fixed MPS bond dimension $\chi$: (a) RQC-1D with $N=101$, (b) PQC-1D with $N=100$, (c) RQC-2D with $L_{x}=L_{y}=12$ (i.e., $N=144$), and (d) PQC-2D with $L_{x}=L_{y}=12$ (i.e., $N=144$). The open diamonds (squares) represent the averaged results of the wavefunction norm $n^{*}$ ($n$) over 10 simulations of the same quantum circuit but with different sets of random parameters, evaluated with (without) the wavefunction norm stabilization procedure. Black dashed lines indicate the norm equal to 1. Here we set $g=0$, i.e., no PtSU since it does not affect the wavefunction norm. Note that the results obtained using the pTEBD algorithm with the wavefunction norm stabilization procedure are almost completely overlapped to each other in this scale.
@@ IV.2.1 Accuracy
First of all, we assess the simulation accuracy of the pTEBD algorithm and demonstrate its ability to achieve precision comparable to that of the sequential MPS algorithm. Here, the simulation precision is determined by the wavefunction fidelity
$\mathcal{F}=|\langle\Psi_{\rm exact}|\Psi(\chi)\rangle|^{2}~{},$ (38)
where $|\Psi_{\rm exact}\rangle$ represents the state obtained by the exact simulation and $|\Psi(\chi)\rangle$ is the state obtained using an MPS with the fixed bond dimension $\chi$. We perform the simulations for quantum circuits with up to 100 (60) physical circuit layers, i.e., before recompiling a quantum circuit to fit an MPS 1D path, in the 1D (2D) cases. The definitions of the physical and compiled circuit depth can be found in Appendix A. Each simulation is repeated 10 times with different sets of random parameters to evaluate the average fidelity, denoted as ¯$\bar{\mathcal{F}}_{\rm MPS(pTEBD)}$ in the sequential MPS (pTEBD) simulations. These results are summarized in Fig. 4.
The simulations for all these four types of quantum circuits show consistent results: for a given bond dimension $\chi$, the accuracy of the pTEBD simulation is comparable to the sequential MPS simulation. In most cases, we find that ¯$\bar{\mathcal{F}}_{\rm pTEBD}\approx\bar{\mathcal{F}}_{\rm MPS}$ and ¯$\bar{\mathcal{F}}_{\rm pTEBD}$ improves with an increase in the number $g$ of PtSU steps, while we observe that the PtSU steps have the opposite effect in some rare cases. Moreover, in cases with relatively small $\chi$ (for example, $\chi=16$ and 64), we notice that ¯$\bar{\mathcal{F}}_{\rm pTEBD}$ significantly increases with applying more PtSU steps. Since $\mathcal{F}$ is highly sensitive to the random parameters used in each quantum gate in the quantum circuit, we also shown its minimum and maximum among the 10 simulations with different sets of random parameters. These extremes obtained in the sequential MPS simulations and the pTEBD simulations are consistent with each other. In some cases [see Fig. 4(b)], the maximum value obtained in the pTEBD simulation can be larger than that obtained in the sequential MPS simulation. To present more clearly the fidelities obtained using these two algorithms, we also show the results for a quantum circuit with the same single set of random parameters in Appendix B.
@@
Figure 6: Elapsed time per circuit layer (averaged over 10 simulations with different sets of random parameters) versus the system size $N$ ($L_{x}=L_{y}$ in 2D cases) obtained using the sequential MPS algorithm (open circles) and the pTEBD algorithm (crosses) with a fixed MPS bond dimension $\chi$: (a) RQC-1D with 300 physical circuit layers, (b) PQC-1D with 300 physical circuit layers, (c) RQC-2D with 100 physical circuit layers, and (d) PQC-2D with 100 physical circuit layers. Note that the number of circuit layers in (c) and (d) is the total number of layers of the quantum circuit after recompiling it to fit an MPS 1D path, i.e., the compiled circuit depth (for more details, see Appendix A). We set $g=0$ for the pTEBD simulations since the performance showing the almost perfect weak scaling does not depend on $g$.
@@ IV.2.2 Numerical stability
IV.2.3 Performance and week scaling
Finally, let us assess the performance of the pTEBD algorithm from two perspectives. On the one hand, we examine the simulation elapsed time per circuit layer, which is defined as the total simulation elapsed time divided by compiled circuit depth (for more details, see Appendix A), versus the system size $N$, evaluating its performance in weak scaling. We always distribute tensors of each 4 adjacent sites in an MPS on one computational node. When $N$ is not divisible by 4, the remaining tensors are distributed on the last node. On the other hand, we compare the elapsed time required to achieve a given simulation precision, measured by wavefunction fidelity $\mathcal{F}$, in both sequential MPS and pTEBD simulations.
As illustrated in Fig. 6, for a fixed bond dimension $\chi$, the pTEBD algorithm consistently exhibits nearly constant simulation times for quantum circuits with various system sizes, achieving excellent weak scaling performance. This is attributed to the absence of any real-space sequential procedures and the lack of a need for global data communication in the pTEBD algorithm [also see Fig. 3(d)]. In contrast, the elapsed time per circuit layer in the sequential MPS algorithm increases linearly with the system size due to its unavoidable sweep procedure. We also expect good strong scaling since, compared with the previous approach in Ref. [28], the IPMC does not increase the computational complexity.
Figure 7: Wavefunction fidelity $\mathcal{F}$ versus elapsed time obtained using the sequential MPS algorithm (open circles) and the pTEBD algorithm (crosses) with a fixed physical circuit depth $D$ (for the definition, see Appendix A): (a) RQC-1D with $N=25$, (b) PQC-1D with $N=24$, (c) RQC-2D with $L_{x}=L_{y}=5$ (i.e., $N=25$), and (d) PQC-2D with $L_{x}=4$ and $L_{y}=6$ (i.e., $N=24$). These results are averaged over 10 simulations of the same quantum circuit but with different sets of random parameters. We set $g=0$ for the pTEBD simulations. Note that the elapsed time increases simply because of the increase of bond dimensions.
It might be more insightful to compare the performance between these two algorithms by studying their elapsed time to achieve the same simulation precision, a measure similarly adopted in a recent study of a 2D tensor network algorithm [47]. This measure demonstrates the practical speedup of the pTEBD algorithm over its sequential counterpart. For this purpose, we reinterpret the fidelity results shown in Fig. 4 and replot these results as a function of the elapsed time of the simulation in Fig. 7. Remarkably, in all cases, the pTEBD simulation reaches the same $\mathcal{F}$ in polynomially shorter time than the corresponding sequential MPS simulation (as also expected from Fig. 6), indicating its efficient utilization of the parallel computing environment.
@@ IV.2.3 Performance and week scaling
In summary, we have proposed an improved parallel MPS compression method that can accurately compress the dimensions of all the virtual bonds in a constant time, regardless of the system size. Simultaneously, it stabilizes the wavefunction norm of the MPS, converging to a value around 1 that is bounded from both sides. Although both the accuracy and the norm stabilization are mathematically proved under the assumption of the canonical form, we have demonstrated its feasibility in simulating unitary quantum dynamics, where the canonical form slightly deviates globally. Moreover, we have numerically shown that the deviated canonical form resulting from the parallel truncation procedure in the parallel MPS compression can be gradually restored through the subsequent PtSU steps. Additionally, we have provided a proof that at most $\frac{N}{2}$ PtSU steps can drive any MPS to its canonical form.
Utilizing the IPMC method, we have proposed a fully real-space parallelizable pTEBD algorithm for efficiently simulating unitary quantum dynamics. Furthermore, we systematically benchmarked the pTEBD algorithm by simulating typical 1D and 2D quantum circuits. Our results demonstrate that the pTEBD algorithm achieves nearly perfect weak scaling performance even for very deep quantum circuits with hundreds of circuit layers. Additionally, it achieves the same simulation precision in polynomially shorter time compared to the sequential MPS algorithm.
@@
As we have demonstrated, the pTEBD algorithm exhibits the same simulation capability as other sequential MPS algorithms. However, due to the inherent limitation of MPS expressibility, simulating quantum circuits with higher-dimensional structures and/or rapidly accumulating large entanglement remains challenging [48, 2, 9]. Therefore, it is of great interest to extend the present parallel approach to recently developed approximate contraction algorithms for more complex tensor networks [49, 50] or to integrate the pTEBD algorithm with other levels of parallelization, such as those utilized in quantum chemistry simulations [51], which are feasible on current supercomputers.
@@ Acknowledgements.
We acknowledge the valuable discussions with Hong-Hao Tu, Kazuma Nagao, Hidehiko Kohshiro, and Shigetoshi Sota. Numerical simulations were performed using the high-performance tensor computing library GraceQ/tensor [52] on the Supercomputer Fugaku installed at RIKEN Center for Computational Science (Project ID No. hp220217 and No. hp230293). This work is supported by Grant-in-Aid for Scientific Research (A) (No. JP21H04446) and Grant-in-Aid for Scientific Research (C) (No. JP22K03479) from MEXT, Japan, as well as JST COI-NEXT (Grant No. JPMJPF2221). Additionally, this work receives support from the Program for Promoting Research on the Supercomputer Fugaku (No. JPMXP1020230411) from MEXT, Japan, and from the COE research grant in computational science from Hyogo Prefecture and Kobe City through the Foundation for Computational Science. This work is also supported in part by the New Energy and Industrial Technology Development Organization (NEDO) (No. JPNP20017).
Appendix A Details of the quantum circuits adopted in the benchmarking simulations
To benchmark the pTEBD algorithm, we select random quantum circuits (RQCs) and parametrized quantum circuits (PQCs), the latter being commonly used in variational quantum algorithms, as the simulation tasks. RQCs have been widely employed in benchmarking NISQ devices [2, 53] and quantum computing simulations [19]. PQCs, being more featured circuits, can reflect the practical performance of devices and simulations for more realistic computational tasks. Therefore, these benchmarks allow for a thorough evaluation of the pTEBD algorithm’s performance in the most realistic and representative cases. In this Appendix, we provide explanations for the construction of these quantum circuits. More detailed features and applications of these quantum circuits can be found in the corresponding references cited below.
A.1 Physical circuit depth vs. compiled circuit depth